A cricketer can throw a ball to a maximum horizontal distance of 100 m. With the same speed how high above the ground can the cricketer throw the same ball?
1
Expert's answer
2018-09-14T10:02:09-0400
The maximum horizontal distance of projectile:
R=v^2/g.
The maximum height of projectile:
H=v^2/2g=R/2=100/2=50 m.